Solve the equation for x: 6.5(4 + x) = 7.5x + 10.
Please help by providing steps. Thanks.

Answer by JBarnum(2146)

6.5(4 + x) = 7.5x + 10
Distribute the 6.5
26+6.5x=7.5x+10
Subtract 6.5x from both sides to get x on just one side of equal sign
26+(6.5x-6.5x)=(7.5x-6.5x)+10
26=1x+10
Subtract 10 from each side
(26-10)=x+(10-10)
16=x
Hope that helps